Kate quit her job, where she earned $120,000 per year, to start her own financial consulting firm. She invested $80,000 of her own funds in furniture, computers, and other assets. During the first year of operation, the firm’s costs were $80,000 for rent on the office building, $420,000 for wages and salaries of employees, and $11,000 for supplies and utilities. The market value of the firm’s assets at the end of the year was $60,000. During the year, the firm billed its clients for 2,700 hours at $250 per hour. The typical rate of return on financial investments in the economy was 8%.

a) in the scenario above, what was Kate's explicit cost ($)?

b) In the scenario above, what was Kate’s cost of capital ($)?

c) In the scenario above, what was Kate’s economic profit ($)?

d) In the scenario above, if Kate earned $130,000 per year at her previous job, what would be her economic profit ($)?
Answer #1

a. Explicit cost = Rent + wages + supplies = 80000 + 420000 + 11000 = 511000

b. Cost of capital = interest forgone on invested amount = 8% * 80000 = 6400

c. Economic profit = Revenue - explicit cost - implicit cost

= 2700 * 250 - 511000 - 6400 - 120000

= 37600

d. Economic profit = Revenue - explicit cost - implicit cost

= 2700 * 250 - 511000 - 6400 - 130000

= 27600